40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3709     Entry time: Wed Oct 13 21:08:40 2010
Author: kiwamu 
Type: Update 
Category: PSL 
Subject: NPRO is still alive 

 The NPRO at the PSL table still can generate 2W laser !  He is still alive.

  When I reduced the temperature to  25 deg, the output power increased to 2W successfully.

  As Steve wrote down in his last entry (see here), the NPRO output was at 1.6 W currently, which is supposed to be 2W.

We were suspicious about the laser crystal's temperature because the current temperature looks a bit high.

In fact the setpoint of the temperature was 45.9 deg instead of 25 deg that is the previous setpoint.

 

ELOG V3.1.3-